Fermi National Laboratory


Accelerator Update for October 26, 2005 - October 28, 2005

Wednesday October 26
The day shift began with Operations monitoring the Recycler stash, with the Tevatron (TeV) in access, with the Antiproton Source (Pbar) stacking, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) in access, and with NuMI, MiniBooNE, and the Meson MCenter experiments taking beam.

Operations began preparing the Recycler at 9:52 AM to stash antiprotons. Extra time was taken to cooldown the existing antiprotons before adding more. The stash ended and Pbar resumed stacking at 11:05 AM.

MCenter resumed taking beam at 11:23 AM.

At 12:42 PM, the Main Injector (MI) P1 transfer line suffered a vacuum cascade. Operators contained the cascade and reset the vacuum pumps.

The TeV access for maintenance work ended at 1:02 PM. Technicians began TeV safety system tests.

Booster experts began conducting high intensity studies at 1:18 PM. These studies interfered with all downstream beam.

A MI expert began a RF study at 1:43 PM.

Experts reported that the P2 transfer line BPM upgrade was completed at 2 PM.

At 5:36 PM, Operations reported that a Pbar Debuncher kicker (D:IKIK) was having problems. An operator called in an expert.

The TeV safety system tests finished at 6:24 PM.

A kicker expert increased the gas pressure of D:IKIK at 6:56 PM. The kicker resumed normal operations.

Operations put the TeV through a dry squeeze at 10:51 PM.

Operations put the TeV through a wet squeeze at 11:52 PM.

Thursday October 27
The midnight shift began with Operations monitoring the stash, with TeV in a wet squeeze, with Pbar stacking, and with MiniBooNE, NuMI, and MCenter taking beam.

Operations put the TeV into shot setup at 12:39 AM.

Operations established store 4473 at 2:52 AM with an initial luminosity of 158.223E30. This store set a New Record Luminosity. Every Fermi luminosity record also sets a New World Luminosity Record. Pbar and MCenter resumed normal operations.

Recycler began preparing at 3:31 AM to stash. The stash ended and Pbar and MCenter resumed normal operations at 4:33 AM.

At 7:53 AM, a Linac RF station (LRF2) tripped off. It reset okay.

Recycler began preparing at 8:09 AM to stash. The Recycler stash ended and Pbar and MCenter resumed normal operations at some later time, probably around 9 AM.

At 8:29 AM, an expert reported that a P2 transfer line power supply (HVF11) had a bad SCR. Power supply techs began repairs around 9:46 AM and finished about 2.5 hours later.

MI and Pbar experts began a bunch rotation study at 1:41 PM. The study ended at 4:22 PM.

Recycler began preparing at 5:23 PM to stash. First, the stash was delayed due to a Booster problem and then it was delayed due to a Pbar problem. The stash ended and Pbar and MCenter resumed normal operations at 7:07 PM.

Recycler began preparing at 11:40 PM to stash.

Friday October 28
The midnight shift began with Operations monitoring store 4473 and the stash, with Pbar stacking, and with MiniBooNE, NuMI, and MCenter taking beam.

The Recycler stash ended and Pbar and MCenter resumed normal operations at 12:40 AM.

At 4:59 AM, Operations reported that the Pelletron tripped off. Operations called in an expert.

The Plans for Friday and the Weekend
The plans for today are to allow TeV experts to conduct an end-of-store study at 10 AM. The rest of the day and the weekend will consist of stacking, stashing, and storing.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 10/28/2005   email Fermilab